OpenStudy (anonymous):

The mean of thirty numbers is 25. If 15 is subtracted from each of the first 12 numbers and 15 is added to each of the last 18 numbers, then what will be the mean of the new set of numbers?

OpenStudy (shubhamsrg):

Whats your attempt ?

Directrix (directrix):

@msingh If the mean of the 25 numbers is 30, then the sum of the 25 numbers is 25*30 = 650. The new sum of the 25 numbers is the following: 650 - 15*12 + 15*18. To get the new mean: Divide (650 - 15*12 + 15*18) by 25.

OpenStudy (anonymous):

@Directrix I think u have made a mistake somewhere See the solution

OpenStudy (anonymous):

The mean of 30 numbers = 25 Then the sum of the 30 numbers =25×30=750 The new sum of the 30 numbers =750-(15×12)+(15×18)

OpenStudy (anonymous):

Thus,the mean of the new set of numbers=[750-(15×12)+(15×18)]/25=33.6

Directrix (directrix):

@msingh Thanks for catching my arithmetic error.